OpenStudy (anonymous):

Emily purchased 6 tokens for an amusement park ride. Each token is worth $0.75. She wants to ride the roller coaster which costs $3.00 a ride. What fractional part of her tokens will she have to use for the ride?(

OpenStudy (anonymous):

please help:(

OpenStudy (anonymous):

.75*3 = 4.50 (total amount of money) this amount is actually 450 cents (4.50*100) $3.00 is 300 cents (3.00*100). 300/450 = 2/3

OpenStudy (anonymous):

does this help ?

OpenStudy (anonymous):

but she has 6 tokens and each one is worth $0.75 so wouldnt that be \[0.75\times6=10.50\]

OpenStudy (anonymous):

no - one second let me write that part for you

OpenStudy (anonymous):

ok!

OpenStudy (anonymous):